# No Other Representation Component Is Needed: Diffusion Transformers Can Provide Representation Guidance by Themselves

### v4

@@ 3.2 Component-Wise Analysis
Effect of projection head for alignment. We finally examine the effect of the projection head for alignment. Surprisingly, We observe that using this simple head to post-possess the student’s output is much better than directly using it to align. We hypothesize this slight operation enables the model to learn more effective hidden representations for subsequent projection head to conduct transformation for final alignment, rather than explicitly aligning the entire latent feature that could potentially disrupt the original generation field that each layer and timestep is responsible for [83, 82, 25]. Hence, we keep using the projection head in future experiments.
@@ 3.3 System-Level Comparison
First, we compare the FID values between vanilla DiT or SiT and the same models trained with SRA. As shown in Figure. 4, diffusion transformers trained with SRA demonstrate substantial improvements in performance at each training step across different types , as well as various sizes. Moreover, similar to the observation in some SSL works [61, 23], we notice that the effect of SRA in a larger size model is more significant, which is probably because the larger model tends to provide richer guidance. It is also worth noting that the benefits of SRA do not saturate even when the models have already achieved a low FID score. We assume that this is likely due to the teacher’s constantly improving capacity throughout training, which allows it to provide better and better representation guidance for the student when the training goes on.
Figure 4: FID comparisons with vanilla DiTs and SiTs across different model sizes on ImageNet 256$\times$256 without classifier-free guidance (CFG).
Table 2: System-level comparison on ImageNet 256$\times$256 with Classifier-free Guidance (CFG). The best and second-best results on each metric are highlighted in bold and underlined.
@@
Figure 5: Selected samples on ImageNet 256$\times$256 from the SiT-XL + SRA. We use classifier-free guidance with $w$ = 4.0. More uncurated samples are provided in Appendix J.
Finally, we provide a quantitative comparison between SiT-XL with SRA and other recent diffusion model methods using classifier-free guidance [34]. Our method already outperforms the original SiT-XL model with 1000 fewer epochs, and it is further improved with longer training. At 800 epochs, SiT-XL with SRA achieves FID of 1.58 and IS of 311.4. It is worth noting that this result is far superior to methods (e.g., MaskDiT [84]) that rely on auxiliary and complex representation training frameworks and is comparable with methods (e.g., REPA [81]) that depend heavily on an external pre-trained representation model. Moreover, because of the progressively higher-quality guidance teacher provide throughout training, we find that our method is much less likely to encounter saturation compared with REPA (we provide detailed results and analysis in Appendix F).
@@ 3.4 Ablation Study
Enhanced representation capacity with SRA. We first compare the representation capacity of vanilla SiT and SiT trained with SRA. As shown in Figure 6a and Figure 6b, SRA consistently improve the quality of latent representation in the diffusion transformer, as indicated by better linear probing results across different blocks and timesteps.
Tight coupling between generation quality and representation guidance in SRA. We then investigate the correlation between generation performance and the representation guidance (detailed experimental setup can be found in Appendix G) in SRA. Figure 6c reveals a strong correlation between linear probing accuracy and FID scores as the teacher network layers for alignment are varied. This finding underscores that the model’s generative capabilities are indeed closely tied to the effectiveness of the self-representation guidance mechanism.
Figure 6: We empirically investigate the effect of representations in SRA. Left and Middle: Linear probing result of vanilla SiT-XL trained for 1400 epochs and SiT-XL + SRA trained for 800 epochs. Right: Linear probing vs. FID plot of SiT-XL + SRA with different teacher’s output layers for alignment (similar plot of DiT + SRA is provided in Appendix H).
@@ 4 Related Work
Here, we highlight key related studies and defer a discussion of other relevant studies to Appendix I.
@@
Representations guidance for diffusion transformer’s pre-training. Many recent works have attempted to introduce representation guidance in diffusion transformer training. MaskDiT [84] and SD-DiT [86] add MAE’s [29] and IBOT’s [85] training paradigm into the original DiT’s training progress. TREAD [42] design a token routing strategy with MAE loss to speed up the diffusion model’s training. REPA [81] utilizes a large-scale data pre-trained representation model for regulating the diffusion model’s latent feature. VA-VAE [80] and MAETok [9] align the latent distribution of the tokenizer with the external representation foundation model and show this alignment is beneficial to resulting diffusion models. Different from this work, we aim to look for representation guidance in the diffusion model itself and its own training paradigm.
EMA model as teacher. Unlike traditional knowledge distillation [27] that uses a well pre-trained model as the teacher, using EMA model [36] as the teacher can be seen as self-distillation because the weight of the teacher model is obtained by the weighted moving average of the student. This method often needs a feasible pretext task to succeed. For example, MoCo [30, 14] sets the teacher’s output as a momentum queue and uses contrastive learning to guide the student model; DINO [8, 61] feeds two views of images to the teacher, and the trainable student then forces the student’s output distribution to be close to that of the teacher. Our work also shares some similarities, where we set aligning the student model’s latent feature in the earlier layer conditioned on higher noise with that in the later layer conditioned on lower noise of the teacher as our pretext task for training.
5 Conclusion and Discussion
@@
In this study, we show that diffusion transformers can provide representation guidance by themselves to boost generation performance with our proposed SRA, which aligns their latent representation in the earlier layer conditioned on higher noise to that in the later layer conditioned on lower noise to progressively enhance the representation learning without external components.
Meanwhile, a few open questions are worth discussing. First, we have observed that the effect of SRA becomes more pronounced when scaling model size. However, owing to the limitation of computational resources, we are unable to further scale the model size and data. Investigating the scalability of SRA in more complex and resource-intensive scenarios (e.g., text-to-image) will be an exciting future direction. Next, similar to other related works [81, 84, 9, 86], our method is also experiment-driven. Exploring theoretical insights into why learning a good representation is beneficial to generation will also be an exciting future direction.
Considering the simplicity and effectiveness of SRA, we believe it will facilitate more future researches that focus on unveiling the intertwined nature of visual representation and generation.
@@ 6 Acknowledgment
Our SRA manuscript referred to some of the color schemes and structural options provided in the REPA [81] manuscript. We sincerely thank REPA’s authors for their fully open-source project. We also thank Sihyun Yu, Sizhe Dang, and Zanyi Wang for the helpful discussions and suggestions during the progress of SRA project.
@@ Appendix A Investigation of Representations in DiT
We also perform a similar analysis with DiT like those have done in Figure 2a (PCA visualization) and Figure 2b (linear probing), the results are showed in Figure 7. In short, we also observe that the representations in DiT basically lead a process from coarse to fine when increasing block layers and decreasing noise level as SiT’s.
Figure 7: We also empirically investigate the representations in diffusion transformers across different blocks and timesteps with the original DiT-XL/2 checkpoint trained for 7M iterations. Similar to SiT, the latent representations in DiT basically follow the coarse to fine process, as block layers increase and noise level decreased.
@@ B.1 Denoising Diffusion Probabilistic Models (DiT)
Denoise-based models [32, 60] aim to model the target distribution ( ${\mathbf{x}}\sim p({\mathbf{x}})$ ) by learning a gradual denoising process that transforms a Gaussian distribution ( $\mathcal{N}(\mathbf{0},\mathbf{I})$ ) into $p({\mathbf{x}})$. Formally, diffusion models learn a reverse process ( $p({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t})$ ) corresponding to a pre-defined forward process ( $q({\mathbf{x}}_{t}|{\mathbf{x}}_{0})$ ), which incrementally adds Gaussian noise to the data starting from $p({\mathbf{x}})$ over a sequence of time steps ( $t\in{1,\dots,T}$ ), with $T>0$ fixed.
For a given ${\mathbf{x}}_{0}\sim p({\mathbf{x}}_{0})$, the forward process ( $q({\mathbf{x}}_{t}|{\mathbf{x}}_{t-1})$ ) is defined as: $q({\mathbf{x}}_{t}|{\mathbf{x}}_{t-1})=\mathcal{N}({\mathbf{x}}_{t};\sqrt{1-\beta_{t}}{\mathbf{x}}_{0},\beta_{t}^{2}\mathbf{I})$, where $\beta_{t}\in(0,1)$ are pre-defined small hyperparameters. DDPM [32] formalizes the reverse process $p({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t})$ as:
@@
$\mathcal{L}_{\text{simple}}=\mathbb{E}_{{\mathbf{x}}_{t},\bm{\epsilon},t}\Big{[}||\bm{\epsilon}-\bm{\epsilon}_{\bm{\zeta}}({\mathbf{x}}_{t},t)||_{2}^{2}\Big{]}.$ (8)
For the covariance term ( $\mathbf{\Sigma}_{\bm{\zeta}}({\mathbf{x}}_{t},t)$ ), DDPM[32] demonstrated that setting it as ( $\sigma_{t}^{2}\mathbf{I}$ ) with ( $\beta_{t}=\sigma_{t}^{2}$ ) is sufficient. Subsequently, Improved-DDPM[60] showed that performance can be enhanced by jointly learning ( $\mathbf{\Sigma}_{\bm{\zeta}}({\mathbf{x}}_{t},t)$ ) along with ( $\bm{\epsilon}_{\bm{\zeta}}({\mathbf{x}}_{t},t)$ ) in a dimension-wise manner using the following objective:
@@
With a sufficiently large $T$ and an appropriate scheduling of $\beta_{t}$, the distribution $p({\mathbf{x}}_{T})$ approaches an isotropic Gaussian distribution. Thus, sampling is achieved by starting from random Gaussian noise and iteratively applying the reverse process ( $p({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t})$ ) to recover a data sample ${\mathbf{x}}_{0}$ [32].
@@ B.2 Stochastic Interpolants Models (SiT)
Unlike DDPMs, flow-based models [51, 49] describe a continuous time-dependent process involving data ( ${\mathbf{x}}_{\ast}\sim p({\mathbf{x}})$ ) and Gaussian noise ( ${\mathbf{\epsilon}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ ) over $t\in[0,1]$:
@@
$\mathcal{L}_{\text{velocity}}=\mathbb{E}_{{\mathbf{x}}_{0},\bm{\epsilon},t}\Big{[}||{\mathbf{v}}_{\bm{\zeta}}({\mathbf{x}}_{t},t)-\dot{\alpha}_{t}{\mathbf{x}}_{0}-\dot{\sigma}_{t}\bm{\epsilon}||^{2}\Big{]}.$ (12)
@@
As shown in [2], any functions $\alpha_{t}$ and $\sigma_{t}$ that satisfy the following three conditions:
@@
lead to an unbiased interpolation process between ${\mathbf{x}}_{0}$ and ${\mathbf{\epsilon}}$. Example choices include linear interpolants ($\alpha_{t}=1-t$, $\sigma_{t}=t$) or variance-preserving (VP) interpolants ($\alpha_{t}=\cos(\frac{\pi}{2}t)$, $\sigma_{t}=\sin(\frac{\pi}{2}t)$) [56].
An advantage of stochastic interpolants is that the diffusion coefficient ( $w_{t}$ ) can be independently selected during sampling with the reverse SDE, even after training. This capability simplifies the design space compared to score-based diffusion models [39].
@@ Appendix C Hyperparameter and More Implementation Details
Table 3: Defualt hyperparameter setup. Unless other otherwise specified, we use these sets of hyperparameters for different models. In our component-wise analysis experiment, settings are also kept the same except those we point out in Table 3.2.
@@
Input dim. 32$\times$32$\times$4 32$\times$32$\times$4 32$\times$32$\times$4 32$\times$32$\times$4 32$\times$32$\times$4 32$\times$32$\times$4
@@
Alignment blocks $83\xrightarrow{}$ 8 $166\xrightarrow{}$ 16 $208\xrightarrow{}$ 20 $73\xrightarrow{}$ 7 $146\xrightarrow{}$ 14 $168\xrightarrow{}$ 16
@@
Objective smooth-$\ell_{1}$ smooth-$\ell_{1}$ smooth-$\ell_{1}$ smooth-$\ell_{1}$ smooth-$\ell_{1}$ smooth-$\ell_{1}$
@@
$\alpha_{t}$ $1-t$ $1-t$ $1-t$
$\sigma_{t}$ $t$ $t$ $t$
$w_{t}$ $\sigma_{t}$ $\sigma_{t}$ $\sigma_{t}$
@@
More implementation details. We implement our models based on the original DiT and SiT implementation. To speed up training and save GPU memory, we use mixed-precision (fp16) with a gradient clipping and FusedAttention [17] operation for attention computation. We also pre-compute compressed latent vectors from raw pixels via stable diffusion VAE [68] and use these latent vectors. We do not apply any data augmentation, but we find this does not lead to a big difference, as similarly observed in MaskDiT [84] and REPA [81]. We also use stabilityai/sd-vae-ft-ema for encoding images to latent vectors and decoding latent vectors to images. For the projection head used for nonlinear transformation, we use two-layer MLP with SiLU activations [21]. The $\lambda$ is set to 0.04 for DiT and 0.2 for SiT initially and followed the rules: $\lambda=\lambda_{\text{init}}*0.1^{\left(\frac{N_{e}-150}{1000}\right)},$ where $N_{e}>150$ is the number of epoch, to change after 150 epochs for ensuring $\mathcal{L}_{\text{gen}}$ and $\mathcal{L}_{\text{sa}}$ at the same scale.
When using SiT-XL to generate images with classifier-free guidance [34], the guidance interval introduced in the study [43] with the same setting used in REPA [81] is applied, which has been demonstrated to yield a slight performance improvement.
Linear probing. We follow the setup used in REPA [81] and I-DAE [13], and use the code-base of ConvNeXt [53] to conduct the experiment. Specifically, we use AdaptiveAvgPooling and a batch normalization layer to process the output of the latent feature by the model, then train a linear layer for 80 epochs. The batch size is set to 4096 with a cosine decay learning rate scheduler, where the initial learning rate is set to 0.001.

### v5

@@ 3.2 Component-Wise Analysis
Effect of projection head for alignment. We finally examine the effect of the projection head for alignment. Surprisingly, We observe that using this simple head to post-possess the student’s output is much better than directly using it to align. We hypothesize this slight operation enables the model to learn more effective hidden representations for subsequent projection head to conduct transformation for final alignment, rather than explicitly aligning the entire latent feature that could potentially disrupt the original generation field that each layer and timestep is responsible for [lsast, prospect, blora]. Hence, we keep using the projection head in future experiments.
@@ 3.3 System-Level Comparison
SRA accelerate the convergence of models in different types and sizes. As our method requires an additional forward pass of the teacher model. For a fair comparison, we report the FID comparison with the baseline under training time. We also give specific training speed and GPU memory usage in Appendix H, as well as FID vs. training time against REPA in Figure 6. As shown in Figure. 4, diffusion transformers trained with SRA demonstrate substantial improvements in performance at the same training time. Moreover, similar to the observation in some SSL works [dinov2, web-ssl], we notice that the effect of SRA in a larger size model is more significant, which is probably because the larger model tends to provide richer guidance. Meanwhile, the benefits of SRA do not saturate even when the models have already achieved a low FID score. We assume that this is likely due to the teacher’s constantly improving capacity, which allows it to provide better and better representation guidance for the student when the training goes on.
Figure 4: FID comparisons with vanilla DiTs and SiTs on ImageNet $\times256$ without classifier-free guidance (CFG). The training time (h) is tested on a single machine with 8 A100 GPUs.
Figure 5: Selected samples on ImageNet $\times256$ from the SiT-XL + SRA. We use classifier-free guidance with $w$ = 4.0. More uncurated samples are provided in Appendix L.
Table 2: FID and Pickscore [pickscore] comparisons with vanilla MMDiT and REPA on COCO2014.
[…]
ODE, NFE=50, Trained for 150K iter
MMDiT 5.86 20.05
MMDiT + REPA 4.60 20.88
MMDiT + SRA 4.85 21.14
Meanwhile, we also conduct text-to-image experiment following REPA to use COCO2014 [coco2014] and MMDiT [sd3] to further demonstrate the effect of SRA, the results in Table 3.3 shows that SRA can naturally extend to text-to-image generation. Our self-representaion alignment strategy works, as evidenced by our experiments: applying SRA to MMDiT without any hyperparameter tuning improves FID ( from 5.66 to 4.75) and PickScore (from 20.65 to 21.14) over the baseline, and is comparable to REPA.
Table 3: System-level comparison on ImageNet $\times256$ with Classifier-free Guidance (CFG). The best and second-best results on each metric are highlighted in bold and underlined.
@@
Table 4: System-level comparison on ImageNet $\times512$ with Classifier-free Guidance (CFG).
[…]
Pixel diffusion
  00VDM++++ - 02.65 - 278.1 - -
  00Simple diffusion 800 04.28 - 171.0 - -
Latent diffusion, Transformer
  00DiT-XL/2 0600 03.04 5.02 240.8 0.84 0.54
  00SiT-XL/2 0600 02.62 4.18 252.2 0.84 0.57
  00MaskDiT 0600 02.50 5.10 256.3 0.84 0.56
  00SiT + REPA 0200 02.08 4.19 274.6 0.83 0.58
  00SiT + SRA (ours) 0200 02.17 4.15 279.3 0.83 0.59
SRA demonstrates superior or comparable performance compared to other methods. We also provide a quantitative comparison between SiT-XL with SRA and other recent methods shown in Table 3.3 and Table 3.3. In ImageNet $\times256$, Our method already outperforms the original SiT-XL model with 1000 fewer epochs, and it is further improved with longer training. At 800 epochs, SRA achieves FID of 1.58 and IS of 311.4. It is worth noting that this result is far superior to methods (e.g., MaskDiT) that rely on external representation task and is comparable with methods (e.g., REPA) that depend heavily on an external pre-trained representation model.
Figure 6: Training time (h) vs. FID plot without CFG.
[…]
SRA shows a more sustained boost than REPA. As shown in Figure 6, although REPA converges quickly at the beginning of training due to its large-scale, pre-trained representation model for guidance, the performance saturates after about 200 epochs. On the contrary, the progressively higher-quality guidance teacher have throughout training in SRA makes our method to provide a more sustained boost without the need for any external models. This observation of saturation effects and potential detrimental impacts of REPA in later stages has also been corroborated by several contemporary studies [repanotwork, waver2025]. We believe this limitation of REPA can be offset by the sustained enhancement facilitated by SRA.
@@ 3.4 Ablation Study
Enhanced representation capacity with SRA. We first compare the representation capacity of vanilla SiT and SiT trained with SRA. As shown in Figure 7((a)) and Figure 7((b)), SRA consistently improve the quality of latent representation in the diffusion transformer, as indicated by better linear probing results across different blocks and timesteps.
Tight coupling between generation quality and representation guidance in SRA. We then investigate the correlation between generation performance and the representation guidance (detailed experimental setup can be found in Appendix I) in SRA. Figure 7((c)) reveals a strong correlation between linear probing accuracy and FID scores as the teacher network layers for alignment are varied. This finding underscores that the model’s generative capabilities are indeed closely tied to the effectiveness of the self-representation guidance mechanism.
Consistent representation improvement during training in SRA. We finally verify whether the representational capacity of the student and teacher models is continuously improving during the training process and whether the teacher can always give better guidance for the student. As shown in Figure 7((d)), the teacher’s representation quality consistently improves throughout training (from 38.1 at 200K iterations to 54.2 at 800K iterations) and consistently outperforms the student model at each stage. This empirical evidence directly supports our claim that the teacher’s improving capacity provides better representation guidance as training proceeds.
Figure 7: We empirically investigate the effect of representations in SRA. (a) and (b): Linear probing result of vanilla SiT-XL trained for 1400 epochs and SiT-XL + SRA trained for 800 epochs. (c): Linear probing vs. FID plot of SiT-XL + SRA with different teacher’s output layers for alignment (similar plot of DiT + SRA is provided in Appendix J). (d): Linear probing results of layers of the teacher and the student used for alignment during training.
@@ 4 Related Work
Here, we highlight key related studies and defer a discussion of other relevant studies to Appendix K.
@@
Representations guidance for diffusion transformer’s training. Many recent works have attempted to introduce representation guidance in diffusion transformer training. MaskDiT [maskdit] and SD-DiT [sddit] add MAE’s [mae] and IBOT’s [ibot] learning task into the original DiT’s training progress. TREAD [tread] design a token routing strategy with MAE loss to speed up the diffusion model’s training. REPA [repa] utilizes a large-scale data pre-trained representation model for regulating the diffusion model’s latent feature. VA-VAE [lightningdit] and MAETok [maetok] align the latent distribution of the tokenizer with the external representation foundation model and show this alignment is beneficial to resulting diffusion models. Different from this work, we aim to look for representation guidance in the diffusion model itself and its own training paradigm.
5 Limitation and Future Work
Due to limitations in compuation resources, we are unable to conduct large-scale text-to-video (t2v) pretraining with SRA. However, we maintain that applying SRA to video domain is conceptually sound for the following reasons. First, even within the image domain (already exist strong encoder like DINOv2), our method achieves performance comparable to that of REPA, this supports our confidence in the effectiveness of approach in text-to-video generation, where currently there is no well-pretrained strong encoder for open-domain videos (there is some video pretrained model, e.g., VideoMAE [videomae], which is still not strong enough for open-domain video encoder, compared to DINOv2 for image encoderAs mentioned by Waver [waver2025], a video generation model uses REPA by employing Qwen2.5-VL [qwen25vl] to extract video representations. It is observed that the early training process benefits from the alignment and the alignment does not bring gain when the training converges. This might stem from the representations extracted from Qwen2.5-VL are not meaningful enough for video generation. Using internal representations of diffusion transformer potentially remedies this..). Second, in the video domain, numerous studies [veo3, vdit] have shown that large-scale pretrained t2v models already capture rich and transferable representations suitable for various understanding tasks, further reinforcing the potential of SRA in video-related applications. Moreover, similar to other related works like REPA, our method is also experiment-driven. Exploring theoretical insights into why learning a good representation is beneficial to generation will also be an exciting future direction.
6 Conclusion
@@
In this study, we show that diffusion transformers can provide representation guidance by themselves to boost generation performance with our proposed SRA, which aligns their latent representation in the earlier layer conditioned on higher noise to that in the later layer conditioned on lower noise to progressively enhance the representation learning without external components. Considering the simplicity and effectiveness of SRA, we believe it will facilitate more future research to extend SRA in other scenarios.
@@ 7 Acknowledgment
Our SRA manuscript referred to some of the color schemes and structural options provided in the REPA [repa] manuscript. We sincerely thank REPA’s authors for their fully open-source project. We also thank Sihyun Yu, Sizhe Dang, and Zanyi Wang for the helpful discussions and suggestions during the progress of SRA project.
@@ Appendix A Investigation of Representations in DiT
We also perform a similar analysis with DiT like those have done in Figure 2(left) (PCA visualization) and Figure 2(right) (linear probing), the results are showed in Figure 8. In short, we also observe that the representations in DiT basically lead a process from coarse to fine when increasing block layers and decreasing noise level as SiT’s.
Figure 8: We also empirically investigate the representations in diffusion transformers across different blocks and timesteps with the original DiT-XL/2 checkpoint trained for 7M iterations. Similar to SiT, the latent representations in DiT basically follow the bad to good process, as block layers increase and noise level decreased.
@@ B.1 Denoising Diffusion Probabilistic Models (DiT)
Denoise-based models [ddpm, improved-ddpm] aim to model the target distribution ( ${\mathbf{x}}\sim p({\mathbf{x}})$ ) by learning a gradual denoising process that transforms a Gaussian distribution ( $\mathcal{N}(\mathbf{0},\mathbf{I})$ ) into $p({\mathbf{x}})$. Formally, diffusion models learn a reverse process ( $p({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t})$ ) corresponding to a pre-defined forward process ( $q({\mathbf{x}}_{t}|{\mathbf{x}}_{0})$ ), which incrementally adds Gaussian noise to the data starting from $p({\mathbf{x}})$ over a sequence of time steps ( $t\in{1,\dots,T}$ ), with $T>0$ fixed.
For a given ${\mathbf{x}}_{0}\sim p({\mathbf{x}}_{0})$, the forward process ( $q({\mathbf{x}}_{t}|{\mathbf{x}}_{t-1})$ ) is defined as: $q({\mathbf{x}}_{t}|{\mathbf{x}}_{t-1})=\mathcal{N}({\mathbf{x}}_{t};\sqrt{1-\beta_{t}}{\mathbf{x}}_{0},\beta_{t}^{2}\mathbf{I})$, where $\beta_{t}\in(0,1)$ are pre-defined small hyperparameters. DDPM [ddpm] formalizes the reverse process $p({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t})$ as:
@@
[…]
For the covariance term ( $\mathbf{\Sigma}_{\bm{\zeta}}({\mathbf{x}}_{t},t)$ ), DDPM[ddpm] demonstrated that setting it as ( $\sigma_{t}^{2}\mathbf{I}$ ) with ( $\beta_{t}=\sigma_{t}^{2}$ ) is sufficient. Subsequently, Improved-DDPM[improved-ddpm] showed that performance can be enhanced by jointly learning ( $\mathbf{\Sigma}_{\bm{\zeta}}({\mathbf{x}}_{t},t)$ ) along with ( $\bm{\epsilon}_{\bm{\zeta}}({\mathbf{x}}_{t},t)$ ) in a dimension-wise manner using the following objective:
@@
With a sufficiently large $T$ and an appropriate scheduling of $\beta_{t}$, the distribution $p({\mathbf{x}}_{T})$ approaches an isotropic Gaussian distribution. Thus, sampling is achieved by starting from random Gaussian noise and iteratively applying the reverse process ( $p({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t})$ ) to recover a data sample ${\mathbf{x}}_{0}$ [ddpm].
@@ B.2 Stochastic Interpolants Models (SiT)
Unlike DDPMs, flow-based models [rectified-flow, flow-matching] describe a continuous time-dependent process involving data ( ${\mathbf{x}}_{\ast}\sim p({\mathbf{x}})$ ) and Gaussian noise ( ${\mathbf{\epsilon}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ ) over $t\in[0,1]$:
@@
[…]
@@
As shown in [si], any functions $\alpha_{t}$ and $\sigma_{t}$ that satisfy the following three conditions:
@@
lead to an unbiased interpolation process between ${\mathbf{x}}_{0}$ and $\mathbf{\epsilon}$}. Example choices include linear interpolants ($\alpha_{t}=1-t$, $\sigma_{t}=t$) or variance-preserving (VP) interpolants ($\alpha_{t}=\cos(\frac{\pi}{2}t)$, $\sigma_{t}=\sin(\frac{\pi}{2}t)$) [sit].
An advantage of stochastic interpolants is that the diffusion coefficient ( $w_{t}$ ) can be independently selected during sampling with the reverse SDE, even after training. This capability simplifies the design space compared to score-based diffusion models [edm].
@@ Appendix C Hyperparameter and More Implementation Details
Table 5: Default hyperparameter setup. Unless other otherwise specified, we use these sets of hyperparameters for different models. In our component-wise analysis experiment, settings are also kept the same except those we point out in Table 3.2.
@@
Input dim. $\times32$$\times$4 $\times32$$\times$4 $\times32$$\times$4 $\times32$$\times$4 $\times32$$\times$4 $\times32$$\times$4
@@
Alignment blocks $3\xrightarrow{}8$ $6\xrightarrow{}16$ $8\xrightarrow{}20$ $3\xrightarrow{}7$ $6\xrightarrow{}14$ $8\xrightarrow{}16$
@@
Objective $\ell_{1} smooth$-$\ell_{1}$ $\ell_{1} smooth$-$\ell_{1}$ $\ell_{1} smooth$-$\ell_{1}$
@@
$\alpha_{t}$ 1t1-t 1t1-t 1t1-t - - -
$\sigma_{t}$ $t$ $t$ $t$ - - -
$w_{t}$ $\sigma_{t}$ $\sigma_{t}$ $\sigma_{t}$ - - -
@@
More implementation details. We implement our models based on the original DiT and SiT implementation. To speed up training and save GPU memory, we use mixed-precision (fp16) with a gradient clipping and FusedAttention [flashatt2] operation for attention computation. We also pre-compute compressed latent vectors from raw pixels via stable diffusion VAE [sd-vae] and use these latent vectors. We do not apply any data augmentation, but we find this does not lead to a big difference, as similarly observed in MaskDiT [maskdit] and REPA [repa]. We also use stabilityai/sd-vae-ft-ema for encoding images to latent vectors and decoding latent vectors to images. For the projection head used for nonlinear transformation, we use two-layer MLP with SiLU activations [silu]. We use smooth-$\ell_{1}$ objective for alignment because we find it perform slightly better than $\ell_{2}$ and its gradient transitions are smoother than $\ell_{1}$. When using SiT-XL to generate images with classifier-free guidance [cfg], the guidance interval introduced in the study [trick_cfg] with the same setting used in REPA [repa] is applied, which has been demonstrated to yield a slight performance improvement.
Sampler. For DiT, we use the DDPM sampler and set the number of function evaluations (NFE) as 250 by default. For SiT, we use the SDE Euler-Maruyama sampler (for SDE with $w_{t}=\sigma_{t}$) and set the NEF as 250 by default. The settings also align with those used in DiT and SiT papers.
Dataset. We use ImageNet [imagenet], where each image is preprocessed to the resolution of $\times256$ or $\times512$ (denoted as ImageNet $\times256$ or ImageNet $\times512$), and follow ADM [adm] for other data preprocessing protocols.
Linear probing. We follow the setup used in REPA [repa] and I-DAE [i-dae], and use the code-base of ConvNeXt [convnext] to conduct the experiment. Specifically, we use AdaptiveAvgPooling and a batch normalization layer to process the output of the latent feature by the model, then train a linear layer for 80 epochs. The batch size is set to 4096 with a cosine decay learning rate scheduler, where the initial learning rate is set to 0.001.